A certain 18-hole golf course has par-3, par-4, and par-5 holes, and there are twice as many par-4 as par-5 holes. How many holes of each type are there if a golfer has a par on every hole for a score of 70?
Answer by stanbon(48568)

A certain 18-hole golf course has par-3, par-4, and par-5 holes, and there are twice as many par-4 as par-5 holes. How many holes of each type are there if a golfer has a par on every hole for a score of 70?
---------------
Let number of par-5 holes be "x"; par-strokes for these is 5x
Number of par-4 holes is "2x"; par-strokes for these is 4(2x)=8x
Number of par-3 holes is "18-x-2x"="18-3x"; par-strokes for these is 3(18-3x)
=54-9x
--------------
EQUATION:
5x+8x+54-9x=70
4x=16
x=4
# of par-5 holes is x=4
# of par-4 holes is 2x=8
# of par-3 holes is 18-3x=6
================
Cheers,
Stan H.